# Gradient-Guided Annealing for Domain Generalization

### v1

@@ 4.4 Sensitivity Analysis
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=5e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to parameter sets that are not close to solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect to training as the search is limited to spaces near the current model parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in the earlier stages of training. The reason for this is that if perturbations are applied near the end of training, the model is thrown far from a local optimum and will need additional iterations to converge.
@@ 5 Conclusions
In this work we researched the properties of gradient conflicts in the context of domain generalization, which led us to the development of GGA, a simple yet effective novel algorithm that aims to identify points in the parameter space where domain gradients align in early stages of training before continuing optimization, ultimately improving model generalization. Through a comprehensive comparative analysis we have demonstrated the efficacy of GGA, as it is able to achieve superior generalization capabilites in standard DG benchmarks and outperform SoTA baselines. What’s more, as GGA is both model and method agnostic it can be utilized by other methods as well. Interestingly, its combination with previous algorithms proves beneficial, as their majority yields improved performance over all benchmarks. Finally, the impact of our algorithm on gradient alignment is validated on additional experiments, where unlike the baseline, domain gradients align during training.
However, our method does not come without some key limitations. First of all, the number of annealing steps and computation of gradient pair similarity adds a computational cost to overall model training. Furthermore, as it is dependent on domain labels, GGA cannot be applied in the single-source DG setup where no information regarding domain distinction is available to the model during training. Additionally, our method is intrinsically connected to the batch size, as larger batch sizes provide increased information towards aligning domain gradients. In future work, we aim to further research the properties of gradient alignment in multi-domain settings and further develop GGA so that the parameter search during the annealing steps may be aware of the model’s current state and not perturb its weights randomly, while also evaluating on additional datasets hindered by domain shift.
@@
  • An extended literature review discussion, helpful for navigating the Domain Generalization literature under the scope of computer vision
  • A complexity analysis regarding the computational overhead of GGA
  • Detailed results for each dataset domain and algorithm, presented in Table 2 of the main text

### v2

@@ 4.4 Sensitivity Analysis
Regarding the magnitude of weight perturbations during the application of GGA, we found that the optimal value was $\rho=5e-5$. As illustrated in Figure 4, a larger magnitude of perturbation led to decreased model performance. Intuitively, the application of larger noise to the model parameters leads to sets that are not close to solution, making it increasingly difficult for the model to converge. On the other hand, smaller perturbations seem to have little to no effect on training as the search is limited to spaces near the current parameters, which is why the model performance falls back close to that of ERM. With regards to the stage of training during which GGA will be applied, we found that the models yielded better performance when the search was initialized in earlier stages.We hypothesize that applying perturbations near the end of training displaces the model from a local optimum, requiring additional iterations to converge.
@@ 5 Conclusions
In this work, we investigated gradient conflicts in the context of domain generalization, leading to the development of GGA; a simple yet effective algorithm that identifies points in the parameter space where domain gradients align early in training before continuing optimization, ultimately enhancing model generalization. Through a comprehensive comparative analysis we have demonstrated the efficacy of GGA, which is able to achieve superior generalization capabilities in standard DG benchmarks and outperform SoTA baselines. What’s more, as GGA is both model and method agnostic it can also be utilized by other methods. Interestingly, its combination with previous algorithms proves beneficial, as their majority yields improved performance over all benchmarks. Finally, we validate the impact of our algorithm on gradient alignment through additional experiments, showing that, unlike the baseline, domain gradients align during training.
However, our method does not come without some key limitations. First of all, The annealing steps and gradient similarity computation add computational overhead. Furthermore, GGA relies on domain labels, making it inapplicable to single-source DG, where domain labels are not available during training. Its effectiveness is also batch-size dependent, as larger batches provide further information regarding gradient alignment. Future work will explore gradient alignment in multi-domain settings and refine GGA to make annealing steps adaptive rather than random, while also evaluating it on datasets affected by domain shift.
Acknowledgement. This work was supported by the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 965231, project REBECCA (REsearch on BrEast Cancer induced chronic conditions supported by Causal Analysis of multi-source data).
@@
  • An extended literature review discussion, helpful for navigating the Domain Generalization literature under the scope of computer vision.
  • A complexity analysis regarding the computational overhead of GGA.
  • Detailed results for each dataset domain and algorithm, presented in Table 2 of the main text.